OpenStudy (anonymous):

If 3a+c=5b and 3a +5b+c=80, what is the value of b?

OpenStudy (anonymous):

8 10 15 20 needs more info.

OpenStudy (diyadiya):

3a+c=5b 3a+5b+c=80 3a+c+5b=80 3a+c can be written as 5b 5b+5b=80 10b=80 divide both sides by 10 so b= ?

OpenStudy (anonymous):

Shouldn't it be -5b?

OpenStudy (diyadiya):

Why ?

OpenStudy (anonymous):

Because it needs to equal to 80, and the 5b needs to get to the other side

OpenStudy (diyadiya):

No 3a + c =5b <-- this is first equation (1) Now 3a+5b+c=80 it can be written as 3a+c+5b=80 (just rearranging ) (3a+c)+5b=80 Now what is 3a+c ?? 3a +c =5b ( from equation (1) ) So substitute it here (5b)+5b=80 10b=80 divide both sides by 10 and b= ?

OpenStudy (anonymous):

8

OpenStudy (diyadiya):

Right :D

OpenStudy (anonymous):

Thanks! :) @Diyadiya

OpenStudy (diyadiya):

Anytime =)